# Consistent text-to-image generation via Scene De-Contextualization

### v1

@@ Abstract
Consistent text-to-image (T2I) generation seeks to produce identity-preserving images of the same subject across diverse scenes, yet it often fails due to a phenomenon called identity (ID) shift. Previous methods have tackled this issue, but typically rely on the unrealistic assumption of knowing all target scenes in advance. This paper reveals that a key source of ID shift is the native correlation between subject and scene context, called scene contextualization, which arises naturally as T2I models fit the training distribution of vast natural images. We formally prove the near-universality of this scene-ID correlation and derive theoretical bounds on its strength. On this basis, we propose a novel, efficient, training-free prompt embedding editing approach, called Scene De-Contextualization (SDeC), that imposes an inversion process of T2I’s built-in scene contextualization. Specifically, it identifies and suppresses the latent scene-ID correlation within the ID prompt’s embedding by quantifying the SVD directional stability to adaptively re-weight the corresponding eigenvalues. Critically, SDeC allows for per-scene use (one scene per prompt) without requiring prior access to all target scenes. This makes it a highly flexible and general solution well-suited to real-world applications where such prior knowledge is often unavailable or varies over time. Experiments demonstrate that SDeC significantly enhances identity preservation while maintaining scene diversity.
@@ 1 Introduction
Methodologically, existing approaches reduce ID shift in line with the paradigm of transfer learning (Tang et al., 2024): Extracting invariance from given heterogeneous data. This requires prior knowledge of the complete target scenes, enabling the generative model to map different scene prompts into corresponding features (Zhou et al., 2024; Liu et al., 2025) or image pseudo labels (Avrahami et al., 2024; Akdemir & Yanardag, 2024) for constructing such a diversified dataset. In practice, however, target scenes are not always available, for instance, in online cases, rendering this assumption unrealistic and limiting the practical applicability of these methods. Critically, the underlying cause of ID shift with T2I models remains largely unclear.
In this work, we consider that the scene plays a context role that would influence the characterization of identity, called scene contextualization (Fig. 1), causing ID shift. This arises because a T2I model is predominantly trained on natural images with a specific data distribution (e.g., caws often appear on the green fields but not in the sea). Consequently, the generated images are constrained to satisfy such internalized priors.
@@
To probe the connection between scene contextualization and ID shift, we formulate a theoretical framework, showing that the contextualization, inherently induced by the attention mechanism, is not only the primary source of ID shift but also inevitable for pre-trained T2I models. Moreover, we derive theoretical bounds on contextualization strength. Building on these insights, we propose a Scene De-Contextualization (SDeC) approach, which effectively realizes the inverse process of scene contextualization. Specifically, SDeC quantifies the directional stability of the subspace spanned by SVD eigenvectors via a forward-and-backward eigenvalue optimization. After that, the latent scene-ID correlation within the ID prompt’s embedding is identified through eigenvalue variations and then suppressed with adaptive eigenvalue weighting. The contextualization-mitigated ID prompt embedding is then reconstructed from the reweighted eigenvalues for subsequent generation. It can work in a one-prompt-per-scene setting, removing reliance on full target scenes.
Our contributions are: (1) We propose a scene contextualization perspective for ID shift with T2I models; (2) We theoretically characterize and quantify this contextualization, leading to a novel SDeC approach for mitigating ID shift per scene without the need for complete target scenes in advance. (3) Extensive experiments show that SDeC can enhance identity preservation, maintain scene diversity, and offer plug-and-play flexibility at per-scene level and across diverse tasks, e.g., integrating pose map and personalized photo, and generative backbones such as PlayGround-v2.5, RealVisXL-V4.0 and Juggernaut-X-V10.
@@ 3 Scene Contextualization
Figure 2: Example of scene contextualization with ID prompt “a photo of a dog” and scene prompt “chasing a frisbee in a park”. Left: Correlation between tokens and the attention similarity matrix reveals that scene tokens influence ID generation (visualization method follows (Hertz et al., 2023)). Right: Similarity between SVD eigenvectors of ID and scene embedding shows they share an overlapping subspace.
@@ 3.2 Bounding the Strength of Contextualization
$0\leq\|T_{\rm sc}\|_{2}\;\leq\;\epsilon\cdot{\|R_{\cap}\|_{2}}\cdot\underbrace{\|U^{\top}W_{V}P_{\cap}\|_{F}}_{\sigma_{\cap}}\;+\;\epsilon\cdot{\|R_{\perp}\|_{2}}\cdot\underbrace{\|W_{V}^{\top}P_{\cap}^{\perp}U\|_{F}}_{\sigma_{\perp}},$ (5)
In Corollary 2, the term $\sigma_{\cap}$ measures the energy shared between ID ($U$) and scene ($\mathcal{Z}_{\mathrm{sc}}^{k}$) subspaces, while $\sigma_{\perp}$ denotes the energy of ID projected into the scene-specific subspace. We consider that the majority of scene-ID interaction takes place via $\sigma_{\cap}$, whilst $\sigma_{\perp}$ allows for holistic coherence. Thus, we only need to minimize $\sigma_{\cap}$.
@@ 4 Scene De-Contextualization
Concretely, we approximate $P_{\cap}$ by identifying the latent scene-ID correlation subspace in $\mathcal{Z}_{\mathrm{id}}$ (Fig. 3(a)), exploiting a learning process. Subsequently, we suppress the correlation subspace (Fig. 3(b)) to reach $\sigma_{\cap}\rightarrow 0$. For clarity, we hereafter denote $\mathcal{Z}_{\mathrm{id}}$ to $\mathcal{Z}_{\mathrm{id}}^{\rm o}$ and in-training ID embedding is denoted as ${\tilde{\mathcal{Z}}}_{\mathrm{id}}$.
@@
We first achieve the directional correlation measurement via an “forward-and-backward” optimization (Fig. 3 (a)): First pulling ${\tilde{\mathcal{Z}}}_{\mathrm{id}}$ closer to the scene $\mathcal{Z}_{\mathrm{sc}}^{k}$ (forward), followed by restoring back to its original position $\mathcal{Z}_{\mathrm{id}}^{\rm o}$ (backward). We start with solving $\mathcal{Z}_{\mathrm{id}}^{\rm o}=U_{\rm id}^{\rm o}\,\Lambda_{\rm id}^{\rm o}V_{\rm id}^{{\rm o}\top}$ by Singular Value Decomposition (SVD) (Stewart, 1993). Let ${\tilde{\mathcal{Z}}}_{\mathrm{id}}=U_{\rm id}^{\rm o}\,\Lambda_{\rm id}V_{\rm id}^{{\rm o}\top}$ with $\Lambda_{\rm id}$ initiated as $\Lambda_{\rm id}^{\rm o}$; We formulate a two-phase optimization problem:
@@
In the two-phase optimization defined in Eq. (6), the forward phase identifies the directions in ${\mathcal{Z}}_{\rm id}$ that align with ${\mathcal{Z}}_{\rm sc}^{k}$, capturing their shared representations. However, some of these directions may also be essential for representing ${\mathcal{Z}}_{\rm id}$ itself. To mitigate potential semantic degradation in ${\mathcal{Z}}_{\rm id}$, we introduce a backward phase that progressively removes these ID-associated components. During this phase, to maximally exclude ID-related information, we set $\beta\gg 0$.
@@
$\Lambda_{\Delta}=|\Lambda^{*}-\Lambda_{\rm id}^{\rm o}|=\mathrm{diag}(v_{1},\dots,v_{i},\dots,v_{r})\;\text{with}\;v_{i}=\;|\lambda_{i}^{*}-\lambda_{i}^{\rm o}|,$ (7)
@@ 5 Experiments
Benchmark Our evaluation uses the ConsiStory+ (Liu et al., 2025), extending the ConsiStory dataset (Tewel et al., 2024) to 192 prompt sets, generating 1292 images with a wider range of subjects, descriptions, and styles.
@@
Competitors We consider two types of competitors. The first is baseline T2I models, including SD1.5 (Rombach et al., 2022a) and SDXL (Podell et al., 2023). The second includes six state-of-the-art consistent T2I methods: BLIP-Diffusion (Li et al., 2023), Textual Inversion (Gal et al., 2022), PhotoMaker (Li et al., 2024), ConsiStory (Tewel et al., 2024), StoryDiffusion (Zhou et al., 2024), and 1Prompt1Story (1P1S) (Liu et al., 2025). The first three are training based, vs training free for the rest and SDeC. For more extensive test, we introduce (1) 1P1S w/o IPCA, with the attention module IPCA removed to focus on its prompt embedding editing; and (2) SDeC+ConsiStory, to test the complementary effect of our method with existing adapter based method ConsiStory. We exclude IP-Adapter (Ye et al., 2023), as its generated characters are homogeneous in pose and layout, with little ability to follow the scene description instruction. The same setting is applied for fair comparison of all compared methods. (see Appendix-E.2).
Table 1: Quantitative comparison. The best and second-best results are marked in bold and underlined, respectively. PE: Prompt embedding Editing; POT: Prompt Operation Time (per image); GenT: Generation Time (per image); BL: BaseLine.
Method
Base
model
PE ID metrics Scenario metrics Infer. time↓ (s)
VRAM↓
Steps
@@
Training
BLIP-Diffusion SD1.5 0.2851 0.8522 0.3957 0.8187 0 1 3.54 26
@@
Training-Free
ConsiStory SDXL 0.2729 0.8604 0.4207 0.8942 0 27 15.58 50
@@ 5.1 Results analysis
Quantitative analysis We draw these observations from Tab. 1: (1) For training-free methods, 1P1S delivers the best result in the ID metrics, whilst suffering serious inter-scene interference (worst DreamSim-B score, also see Fig. 6 in Appendix-F), largely unacceptable for consistent T2I generation. In contrast, our SDeC strikes the best balance between ID consistency and scene diversity. (2) Without the attention IPCA, 1P1S is outperformed by SDeC across all metrics. That indicates that our prompt embedding editing is superior, even without the need for all target scenes in advance. (3) ConsiStory lags behind SDeC in ID metrics, but excels in scene background. (4) SDeC is well complementary with ConsiStory to further push the performance, as they address distinct aspects. (5) Interestingly, training-based methods are even outpaced by most tree-free counterparts in ID consistency, with extra disadvantage in efficiency. (6) In terms of memory and inference time, SDeC introduces negligible overhead on top.
@@
User study
To complement those evaluation metrics, we further conduct a user study. We compare with top alternatives: PhotoMaker, ConsiStory, StoryDiffusion, and 1P1S. Specifically, the test images were generated by 30 random prompt sets from ConsiStory+. A total of 20 volunteers were invited to pick which image best balances among ID consistency, scene diversity, and prompt alignment. We measure the performance using the percentage of wins. Tab. 2 shows that SDeC best matches human preference.
Qualitative analysis For visual comparison, we show a couple of examples in Fig. 4. For the robotic elephant case, ConsiStory presents varying robotic styles, whilst 1P1S suffers with the scene interference. For the cup of hot chocolate case, the ID shift issue becomes more acute with all previous methods. Instead, SDeC still does a favored job. More qualitative results are provided in Appendix-I.
@@ 5.2 Further Analysis
Additionally, relative to SDeC, the $P_{\cap}$ estimated by these two variant approaches is inherently less reliable. When applied to contextualization carrier suppression, it introduces additional and unpredictable interference into the ID embedding. Through the coupling effect of attention, this interference propagates into the scene generation process, ultimately amplifying discrepancies across scenes. This is why the variant approaches have a higher DreamSim-B score than SDeC. The typical qualitative comparison results can be found in Appendix-G.1.
@@
Additionally, our prompt editing design operates through the attention module. Accordingly, the most direct evidence of contextualization control is to examine the correlation between token embeddings and the attention similarity matrix. Fig. 5-Bottom visualizes it as $\mathcal{Z}_{\rm{id}}^{*}$ drives the generation. With bright green denoting high correlation, we observe that, except for image #7, SDeC yields darker regions with more pronounced subject silhouettes than SDXL. These observations suggest that SDeC reduces the scene contextualization and provides an intuitive explanation for SDeC’s effectiveness. The more model analyses are provided in Appendix-H.
Analysis of a proprietary commercial product Google’s latest flagship T2I product, Nano Banana (Google & DeepMind, 2025), has demonstrated impressive ID-preservation ability. While we cannot integrate our method, we design an interesting test to inspect how this system might work, which may reveal additional insights for future work. The results and speculative analysis are presented in Appendix-K.

### v2

@@ Abstract
Consistent text-to-image (T2I) generation seeks to produce identity-preserving images of the same subject across diverse scenes, yet it often fails due to a phenomenon called identity (ID) shift. Previous methods have tackled this issue, but typically rely on the unrealistic assumption of knowing all target scenes in advance. This paper reveals that a key source of ID shift is the native correlation between subject and scene context, called scene contextualization, which arises naturally as T2I models fit the training distribution of vast natural images. We formally prove the near-universality of this scene-ID correlation and derive theoretical bounds on its strength. On this basis, we propose a novel, efficient, training-free prompt embedding editing approach, called Scene De-Contextualization (SDeC), that imposes an inversion process of T2I’s built-in scene contextualization. Specifically, it identifies and suppresses the latent scene-ID correlation within the ID prompt’s embedding by quantifying SVD directional stability to adaptively re-weight the corresponding eigenvalues. Critically, SDeC allows for per-scene use (one scene per prompt) without requiring prior access to all target scenes. This makes it a highly flexible and general solution well-suited to real-world applications where such prior knowledge is often unavailable or varies over time. Experiments show that SDeC markedly enhances identity preservation while maintaining scene diversity.
@@ 1 Introduction
Methodologically, existing approaches reduce ID shift in line with the paradigm of transfer learning (Tang et al., 2024): Extracting invariance from given heterogeneous data. This requires prior knowledge of the complete target scenes, enabling the generative model to map different scene prompts into corresponding features (Zhou et al., 2024; Liu et al., 2025) or image pseudo labels (Avrahami et al., 2024; Akdemir & Yanardag, 2024) for constructing such a diversified dataset. In practice, however, target scenes are not always available [^1] In real-world projects (e.g., films, games, or story creation), the full set of final scenes, their content, and their order are often refined and finalized over numerous iterative changes, making it impossible to know all subsequent scene contexts in advance. Efficiency dictates generating images online based on the current scene description, without the need for repeatedly re-generating all previous scenes (avoiding exponential complexity)., rendering this assumption unrealistic and limiting the practical applicability of these methods. Critically, the underlying cause of ID shift with T2I models remains largely unclear.
In this work, we consider that the scene plays a context role that would influence the characterization of identity, called scene contextualization (Fig. 1), causing ID shift. This arises because a T2I model is predominantly trained on natural images with a specific data distribution (e.g., cows often appear on the green fields but not in the sea). Consequently, the generated images are constrained to satisfy such internalized priors.
@@
To probe the connection between scene contextualization and ID shift, we formulate a theoretical framework, showing that the contextualization, inherently induced by the attention mechanism, is not only the primary source of ID shift but also inevitable for pre-trained T2I models. Moreover, we derive theoretical bounds on contextualization strength. Building on these insights, we propose a Scene De-Contextualization (SDeC) approach, which effectively realizes the inverse process of scene contextualization. Specifically, SDeC quantifies the directional stability of the subspace spanned by Singular Value Decomposition (SVD) (Stewart, 1993) eigenvectors via a forward-and-backward eigenvalue optimization. After that, the latent scene-ID correlation within the ID prompt’s embedding is identified through eigenvalue variations and then suppressed with adaptive eigenvalue weighting. The contextualization-mitigated ID prompt embedding is then reconstructed from the reweighted eigenvalues for subsequent generation. It can work in a one-prompt-per-scene setting, removing reliance on full target scenes.
Our contributions are: (1) We propose a scene contextualization perspective for ID shift with T2I models; (2) We theoretically characterize and quantify this contextualization, leading to a novel SDeC approach for mitigating ID shift per scene without the need for complete target scenes in advance. (3) Extensive experiments show that SDeC can enhance identity preservation, maintain scene diversity, and offer plug-and-play flexibility at per-scene level and across diverse tasks, e.g., integrating pose map and personalized photo, and generative backbones such as PlayGround-v2.5, RealVisXL-V4.0, Juggernaut-X-V10, SD3, and Flux.
@@ 3 Scene Contextualization
Figure 2: Example of scene contextualization with ID prompt “a photo of a dog” and scene prompt “chasing a frisbee in a park”. Left: Correlation between tokens and attention similarity matrix shows that scene tokens affect ID generation (visualization follows (Hertz et al., 2023)). Right: Similarity between SVD eigenvectors of ID and scene embedding shows they share an overlapping subspace.
@@ 3.2 Bounding the Strength of Contextualization
$0\leq\|T_{\rm sc}\|_{2}\;\leq\;\epsilon\cdot{\|R_{\cap}\|_{2}}\cdot\underbrace{\|U^{\top}W_{V}P_{\cap}\|_{F}}_{\sigma_{\cap}}\;+\;\epsilon\cdot{\|R_{\perp}\|_{2}}\cdot\underbrace{\|W_{V}^{\top}P_{\cap}^{\perp}U\|_{F}}_{\sigma_{\perp}}.$ (5)
In Corollary 2, $\sigma_{\cap}$ measures the energy shared between ID ($U$) and scene ($\mathcal{Z}_{\mathrm{sc}}^{k}$) subspaces, while $\sigma_{\perp}$ denotes the energy of ID projected into the scene-specific subspace. We consider that the majority of scene-ID interaction takes place via $\sigma_{\cap}$, whilst $\sigma_{\perp}$ allows for holistic coherence. Thus, we only need to minimize $\sigma_{\cap}$. This decomposition further reveals a balance between lowering scene–ID interaction and maintaining coherence, indicating an inherent disentanglement–coherence trade-off.
@@ 4 Scene De-Contextualization
Concretely, we approximate $P_{\cap}$ by identifying the latent scene-ID correlation subspace in $\mathcal{Z}_{\mathrm{id}}$ (Fig. 3(a)), exploiting a learning process. We then suppress this subspace (Fig. 3(b)) to reach $\sigma_{\cap}\rightarrow 0$. For clarity, we hereafter denote $\mathcal{Z}_{\mathrm{id}}$ to $\mathcal{Z}_{\mathrm{id}}^{\rm o}$ and in-training ID embedding is denoted as ${\tilde{\mathcal{Z}}}_{\mathrm{id}}$.
@@
We first achieve the directional correlation measurement via an “forward-and-backward” optimization (Fig. 3 (a)): First pulling ${\tilde{\mathcal{Z}}}_{\mathrm{id}}$ closer to the scene $\mathcal{Z}_{\mathrm{sc}}^{k}$ (forward), followed by restoring back to its original position $\mathcal{Z}_{\mathrm{id}}^{\rm o}$ (backward). We start with solving $\mathcal{Z}_{\mathrm{id}}^{\rm o}=U_{\rm id}^{\rm o}\,\Lambda_{\rm id}^{\rm o}V_{\rm id}^{{\rm o}\top}$ by SVD. Let ${\tilde{\mathcal{Z}}}_{\mathrm{id}}=U_{\rm id}^{\rm o}\,\Lambda_{\rm id}V_{\rm id}^{{\rm o}\top}$ with $\Lambda_{\rm id}$ initiated as $\Lambda_{\rm id}^{\rm o}$; We formulate a two-phase optimization problem:
@@
In the two-phase optimization defined in Eq. (6), the forward phase identifies the directions in ${\mathcal{Z}}_{\rm id}$ that align with ${\mathcal{Z}}_{\rm sc}^{k}$, capturing their shared representations. However, some of these directions may also be essential for representing ${\mathcal{Z}}_{\rm id}$ itself. To mitigate potential semantic degradation in ${\mathcal{Z}}_{\rm id}$, we introduce a backward phase that progressively recovers these ID-associated components.
@@
$\Lambda_{\Delta}=|\Lambda^{*}-\Lambda_{\rm id}^{\rm o}|=\mathrm{diag}(v_{1},\dots,v_{i},\dots,v_{r})\;\text{with}\;v_{i}=\;|\lambda_{i}^{*}-\lambda_{i}^{\rm o}|.$ (7)
@@ 5 Experiments
Benchmark Our evaluation uses the ConsiStory+ (Liu et al., 2025), extending the ConsiStory dataset (Tewel et al., 2024) to 192 prompt sets, generating 1292 images with a wider range of subjects, descriptions, and styles. The scene includes all contextual factors: not only environmental attributes (e.g., lighting, style, background elements) but also actions, behaviors, or temporary states associated with the subject in that specific image.
@@
Competitors We consider two types of competitors. The first is baseline T2I models, including SD1.5 (Rombach et al., 2022a) and SDXL (Podell et al., 2023). The second includes six state-of-the-art consistent T2I methods: BLIP-Diffusion (Li et al., 2023), Textual Inversion (Gal et al., 2022), PhotoMaker (Li et al., 2024), ConsiStory (Tewel et al., 2024), StoryDiffusion (Zhou et al., 2024), and 1Prompt1Story (1P1S) (Liu et al., 2025). The first three are training based, vs training free for the rest and SDeC. For more extensive test, we introduce (1) 1P1S w/o IPCA, with the attention module IPCA removed to focus on its prompt embedding editing; and (2) SDeC+ConsiStory, to test the complementary effect of our method with existing adapter based method ConsiStory. We exclude IP-Adapter (Ye et al., 2023) from comparison, as its generated characters are homogeneous in pose and layout, with little ability to follow the scene description instruction. The same setting is applied to all compared methods. (see Appendix-E.2).
Table 1: Quantitative comparison. The best and second-best results are marked in bold and underlined, respectively. PE: Prompt embedding Editing; POT: Prompt Operation Time (per image); GenT: Generation Time (per image).
Method Base model PE ID metrics Scenario metrics Infer. time↓ (s) VRAM↓ (GB) Steps
@@
Training BLIP-Diffusion SD1.5 0.2851 0.8522 0.3957 0.8187 0 1 3.54 26
@@
Training-Free ConsiStory SDXL 0.2729 0.8604 0.4207 0.8942 0 27 15.58 50
@@ 5.1 Results analysis
Quantitative analysis We draw these observations from Tab. 1: (1) For training-free methods, 1P1S delivers the best result in the ID metrics, whilst suffering serious inter-scene interference (worst DreamSim-B score, also see Fig. 7 in Appendix-F for the typical qualitative evidence), largely unacceptable for consistent T2I generation. In contrast, our SDeC strikes the best balance between ID consistency and scene diversity. (2) Without the attention IPCA, 1P1S is outperformed by SDeC across all metrics. That indicates that our prompt embedding editing is superior, even without the need for all target scenes in advance. (3) ConsiStory lags behind SDeC in ID metrics, but excels in scene background. (4) SDeC is well complementary with ConsiStory to further push the performance, as they address distinct aspects. (5) Interestingly, training-based methods are even outpaced by most tree-free counterparts in ID consistency, with extra disadvantage in efficiency. (6) In terms of memory and inference time, SDeC introduces negligible overhead on top (see Appendix H.1 for further discussion).
@@
User study To complement those evaluation metrics, we further conduct a user study. We compare with top alternatives: PhotoMaker, ConsiStory, StoryDiffusion, and 1P1S. Specifically, the test images were generated by 30 random prompt sets from ConsiStory+. A total of 20 volunteers were invited to pick which image best balances among ID consistency, scene diversity, and prompt alignment. We measure the performance using the percentage of wins. Tab. 2 shows that SDeC best matches human preference.
Qualitative analysis For visual comparison, we show a couple of examples in Fig. 4. For the robotic elephant case, ConsiStory presents varying robotic styles, whilst 1P1S suffers from scene interference. For the cup of hot chocolate case, the ID shift issue becomes more acute with prior methods. Instead, SDeC still does a favored job. More qualitative results are provided in Appendix-G.
@@ 5.2 Further Analysis
Additionally, relative to SDeC, the $P_{\cap}$ estimated by these two variant approaches is inherently less reliable. When applied to contextualization carrier suppression, it introduces additional and unpredictable interference into the ID embedding. Through the coupling effect of attention, this interference propagates into the scene generation process, ultimately amplifying discrepancies across scenes. This is why the variant approaches have a higher DreamSim-B score than SDeC.
As a supplement to Tab. 3, in Fig. 5, we present a qualitative comparison between methods SDeC w/o soft-estimation, SDeC w/o abs-excursion, and SDeC. It is seen that the qualitative results are consistent with the data in Tab. 3. In particular, method SDeC w/o abs-excursion performs worse than method SDeC w/o soft-estimation, which supports our previous discussion. Specifically, SDeC w/o abs-excursion disregards the inherent importance of different SVD eigen directions, resulting in greater semantic loss and severe subject deformation (for example, see image #2 in Fig. 5-Right). In contrast, SDeC w/o soft-estimation, which employs a matrix transformation approach, just suffers from less accurate in identifying latent overlapping subspaces, but remain the key semantic information. Thus, it achieves better ID preservation than SDeC w/o abs-excursion.
@@
Additionally, our prompt editing design operates through the attention module. Accordingly, the most direct evidence of contextualization control is to examine the correlation between token embeddings and the attention similarity matrix. Fig. 6-Bottom visualizes it as $\mathcal{Z}_{\rm{id}}^{*}$ drives the generation. With bright green denoting high correlation, we observe that, except for image #7, SDeC yields darker regions with more pronounced subject silhouettes than SDXL. These observations suggest that SDeC reduces the scene contextualization and provides an intuitive explanation for SDeC’s effectiveness. The more model analyses are provided in Appendix-H and Appendix-I.
Analysis of a proprietary commercial product Google’s latest flagship T2I product, Nano Banana (Google & DeepMind, 2025), has demonstrated impressive ID-preservation ability. While we cannot integrate our method, we design an interesting test to inspect how this system might work, which may reveal additional insights for future work. The results and speculative analysis are presented in Appendix-J.